Hans paid $1.50 each for programs to the game. He sold all but 20 of them for $3.00 each and made a profit of $15.00. How many programs did he buy?
I tried the equation: 1.50x + 3(x-20)=15.00
I don't think this is right? Can you help me? Thank You.
The ISBN number that I typed is exactly correct.

Found 2 solutions by checkley71, BGluvsmath:
Answer by checkley71(8403)

1.50X+15.00=(X-20)3.00
1.50X+15.00=3X-60.00
1.50X-3X=-60.00-15.00
-1.50X=-75.00
X=-75.00/-1.5
X=50 PROGRAMS WERE BOUGHT & 30 SOLD.
PROOF
50*1.50+15.00=(50-20)3.00
75.00+15.00=30*3.00
90.00=90.00

Answer by BGluvsmath(64)

First, profit is the selling price minus cost. So, I think a better equation would be: 1.5(x-20)=15 (Hans made a profit of $1.50 per program sold, for a total of $15.)
Going on to solving the equation:
1.5(x-20)=15
distribute: 1.5x-30=15
add 30 to both sides: 1.5x=45
divide each side by 1.5: x=30